OpenStudy (anonymous):

Christa buys 14 flowerpots. Some are small and cost $2.40 each. The rest are large and cost $5.60 each. She spends a total of $49.60. Which equation models this situation? Let s represent the number of small flowerpots she buys. A. 5.6s + 2.4(s + 14) = 49.6 B. (2.4 + 9.6)(14 – s) = 49.6 C. 2.4s + 5.6(14 – s) = 49.6 D. 5.6s + 2.8(14) = 49.6

OpenStudy (anonymous):

@charko

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@heavenfun13

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

anyone??!! D;

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

kk as long as your helping ;D

OpenStudy (anonymous):

c seems to make sense

OpenStudy (anonymous):

kk ill tell you if it is right ;D ;3 ;)

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

Marie has $830 in her bank account and withdraws $60 each month. Denise has $970 in her bank account and withdraws $80 each month. In how many months will Marie and Denise have equal amounts of money in their accounts? A. 4 B. 5 C. 7 D. 9

OpenStudy (anonymous):

@charko

OpenStudy (anonymous):

@heavenfun13

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

wassup

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

since you're looking for the time when their money left is equal, then let x be the number of months 830-60x=970-80x now solve for x

OpenStudy (anonymous):

7 months

OpenStudy (anonymous):

7

OpenStudy (anonymous):

everything was right!!!!!!!!!!!!!!!!!!!!!!! ;D every answer yall gave me was all correct and there was 30 questions!!!! ;D yall are my heros

OpenStudy (anonymous):

i got a 100%

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i am so happy

OpenStudy (anonymous):

i havent been able to get a 100 on any math test exept like 80 and 85 and thats it but i got a 100% i am soo happy thank you so much guys

OpenStudy (anonymous):

Any time

OpenStudy (anonymous):

;D

OpenStudy (anonymous):

can i ask you another one?

OpenStudy (anonymous):

@charko

OpenStudy (anonymous):

Marco and Drew stacked boxes on a shelf. Marco lifted 9 boxes and Drew lifted 14 boxes. The boxes that Drew lifted each weighed 8 lb less than the boxes Marco lifted. Which expression represents the total number of pounds Drew lifted? Let m represent the weight of the boxes that Marco lifted.

OpenStudy (anonymous):

A. m − 8 B. 14(m − 8) C. 9(m + 8) D. 112m

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Two stores sell the same refrigerator for the same original price. Store A advertises that the refrigerator is on sale for 15% off the original price. Store B advertises that it is reducing the refrigerator’s price by $150. When Stephanie compares the sale prices of the refrigerator in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the refrigerator’s original price. A. 0.85p = p – 150 B. 0.85(p – 150) = p C. 0.15p = p − 150 D. 0.15p = p + 180

OpenStudy (anonymous):

@charko

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

0.85p = p – 150

OpenStudy (anonymous):

thx i only have 5 of these and next is #3

OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

OpenStudy (anonymous):

55mph

OpenStudy (anonymous):

thx 2 more?

OpenStudy (anonymous):

Gretchen and Ezia received equal scores on a test made up of multiple choice questions and an essay. Gretchen got 18 multiple choice questions correct and received 19 points for her essay. Ezia got 15 multiple choice questions correct and received 31 points for her essay. How many points was each multiple choice question worth

OpenStudy (anonymous):

15x +31=y 18x +19=y 18x+19=15x+31 3x=12 x=4 which is your answer

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

In May, Liam and Charlie had the same amount of money in their savings accounts. In June, Liam deposited $140 into his account. Charlie said he increased the money in his account by 5%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in May

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

amt of money Liam = a Charlie = b a=b a + 140 b + 0.05b a+140 = a + 0.05a 140 = 0.05a 2800 = a

OpenStudy (anonymous):

thank you so much you a life saver

OpenStudy (anonymous):

100%!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! ;D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

u go to k12

OpenStudy (anonymous):

you are my best friend the angle helping and watching over me!!!!!

OpenStudy (anonymous):

m real name is heaven

OpenStudy (anonymous):

really? ;D

OpenStudy (anonymous):

yea lol

OpenStudy (anonymous):

What is 348,000,000,000,000 estimated as the product of a single digit and a power of 10

OpenStudy (anonymous):

well you are my best friend

OpenStudy (anonymous):

3 x 10^14

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

are u in k12

OpenStudy (anonymous):

yeah! ;D

OpenStudy (anonymous):

me too what grade

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

Mr.Montelaro

OpenStudy (anonymous):

i am in 7th

OpenStudy (anonymous):

well im not smart at all so thats why im asking for help xD

OpenStudy (anonymous):

What is 0.0000000734 estimated as the product of a single digit and a power of 10

OpenStudy (anonymous):

single digit means we have to take one digit only which will be 7 7*10^(-8)

OpenStudy (anonymous):

7.34*10^-8

OpenStudy (anonymous):

thx but your sure its in piots? cause what all i have seen is no pionts xD but as long as you help me and have an answer

OpenStudy (anonymous):

Which numbers are greater than 3 × 10−7? Choose all answers that are correct. A. 9 × 10−9 B. 2 × 10−5 C. 6 × 10−7 D. 5 × 10−8 E. 2 × 10−7

OpenStudy (anonymous):

6x10^-7 2x10^-5

OpenStudy (anonymous):

What is the best order-of-magnitude estimate for 29,300

OpenStudy (anonymous):

10^4

OpenStudy (anonymous):

What is the best order-of-magnitude estimate for 0.00006

OpenStudy (anonymous):

i think ill be done after this and if a am thank you so much for your time and your help

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

do you know it?

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@charko

OpenStudy (anonymous):

10-5

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

lulz thx

OpenStudy (anonymous):

C. s = 9, 14-9 = 5

OpenStudy (anonymous):

OH NO I MADE A 60% AND I DIDNT MASTER IT AND I CANT TAKE IT ANYMORE TIMES SO I NEED TO MOVE ON TO ONE MORE BEFORE IM FINISHED TODAY

JoelTheBoss (joel_the_boss):

sorry not the best with math and ^^ is good lol

OpenStudy (anonymous):

8.6e+12 What is the standard form of the number in the calculator’s display? The calculator shows eight point six e plus twelve A. 8,600,000,000,000 B. 860,000,000,000 C. 0.0000000000086 D. 0.00000000000086

OpenStudy (anonymous):

what accures in a year in a week but never in a day? Trick question i know it but do you?

OpenStudy (fateal):

i think you should make different posts n ask more questions???